# Maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$

### v1

@@ Abstract.
We introduce maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ whose deformation space can be understood as the cotangent bundle of Weil-Petersson universal Teichmüller space $T_{0}(1)$. We show that the Mess map is a symplectic diffeomorphism from $T^{*}T_{0}(1)$ to $T_{0}(1)\times T_{0}(1)$ for a canonical symplectic form on $T^{*}T_{0}(1)$ and the difference of two Weil-Petersson symplectic forms on $T_{0}(1)\times T_{0}(1)$. Moreover, we show that the anti-holomorphic energy functional of the induced Gauss map is a Kähler potential of the canonical symplectic form over certain subspaces of $T^{*}T_{0}(1)$.
2010 Mathematics Subject Classification 30F60, 32G15, 53C50, 53A10 .
@@ 1. Introduction
In the seminal work in [15], Takhtajan and Teo introduced a subspace of the universal Teichmüller space $T(1)$, which is equipped with a Hermitian inner product. This space denoted by $T_{0}(1)$ is the Weil-Petersson universal Teichmüller space in the sense that the space $T_{0}(1)$ has rich structures derived from the Weil-Petersson inner product and the associated symplectic 2-form $\omega_{\mathrm{WP}}$. Moreover, in [15], they also introduced the universal Liouville action $S$ for an element in $T_{0}(1)$ and proved that $S$ is a Kähler potential function of $\omega_{\mathrm{WP}}$ such that
(1.1) ¯$\partial\bar{\partial}S=-2\sqrt{-1}\,\omega_{\mathrm{WP}}\qquad\text{over}\quad T_{0}(1).$
Here $\partial$ and $\bar{\partial}$ denote the holomorphic and anti-holomorphic derivatives on $T_{0}(1)$ respectively. The equality (1.1) can be considered as the universal version of the results which have been proved for the classical Liouville actions over Teichmüller spaces of Riemann surfaces in [16], [17], [14], [12].
An element in $T_{0}(1)$ can be represented as a quasi-circle in the complex plane with a certain condition, which is called Weil-Petersson curve. After its introduction in [15], this has drawn many attentions and has become a very active research object in various research subjects. In particular, Wang in [23] introduced the Loewner energy $I_{L}$ for a Weil-Petersson curve using the energy of driving function which naturally appears in the Shramm Loewner equation and showed the following equality
@@
This equality is rather surprising since the same quantity can be defined in terms of totally different constructions developed in Teichmüller theory and Schramm Loewner theory.
In a recent work [3] about the Weil-Petersson curve, Bishop extensively studied various aspects of Weil-Petersson curves in the view points of analysis and geometry. In particular, in [3], Bishop considered an invariant for a minimal surface in the hyperbolic 3-space $\mathbb{H}^{3}$ bounding a given Weil-Petersson curve in ¯$\partial\overline{\mathbb{H}}^{3}$. One drawback of this approach is the non-uniqueness of the bounding minimal surfaces in $\mathbb{H}^{3}$. Hence, it seems to be difficult to relate an invariant of a bounding minimal surface to the universal Liouville action or equivalently the Loewner energy of the Weil-Petersson curve.
Motivated by the works in [15], [23], and [3], we consider a maximal surface in the Anti de Sitter 3-space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, which is a Lorentzian counter part of a minimal surface in $\mathbb{H}^{3}$. More precisely, we introduce a maximal disc of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ such that its boundary is the graph of a quasisymmetric homeomorphism of $\mathbb{S}^{1}$ representing an element in $T_{0}(1)$. By the work [4], such a bounding maximal surface exists uniquely for a given element in $T_{0}(1)$. Allowing a deformation of conformal structures in $T_{0}(1)$ on the given (real) maximal disc, we can see that the cotangent space $T^{*}T_{0}(1)$ parametrizes the embedded maximal discs of Weil-Petersson class in the Anti de Sitter 3-space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. Actually, we do not need the Weil-Petersson condition to apply the result of [4]. But, it turns out that the Weil-Petersson condition is very essential for results in this paper, which we are going to explain.
One of the basic properties for a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is the existence of a pair of harmonic maps $F_{\pm}$ induced by the Gauss map of $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. In general, it is not possible to get a finite valued energy of the harmonic maps $F_{\pm}$ since the domain $\Sigma$ is not compact. But, in Proposition 3.7, it is proved that the Weil-Petersson condition for a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ implies the finiteness of the anti-holomorphic energy of the harmonic maps $F_{\pm}$ on $\Sigma$.
Following the original construction of Mess in his work [10], one can introduce so called Mess map
@@
It has been expected that the map $\mathrm{Mess}$ in (1.3) is a symplectic diffeomorphism. We prove this property for the restriction of $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$. Recall that the Weil-Petersson symplectic form $\omega_{\mathrm{WP}}$ is defined on $T_{0}(1)$ by using the Hilbert manifold structure on $T_{0}(1)$ in [15]. It is also crucial to use the Hilbert manifold structure on $T_{0}(1)$ to introduce the canonical symplectic form $\omega_{\mathrm{C}}$ on $T^{*}T_{0}(1)$. Then, our result for this states as follows
@@
The map $\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$ is a symplectic diffeomorphism, that is,
Here $\mathrm{Mess}_{\pm}:=\pi_{\pm}\circ\mathrm{Mess}$ where $\pi_{\pm}$ denotes the projection map from $T_{0}(1)\times T_{0}(1)$ onto the first or second factor respectively.
A related result to Theorem 1.1 was given in [7] at the origin of $T^{*}T(1)$ and $T(1)\times T(1)$. But, it is not clear which symplectic structures are referred to in [7] considering that a main result of [15] is the construction of the symplectic form $\omega_{\mathrm{WP}}$ on $T_{0}(1)$.
Let us put $T_{0}(1)^{+}$, $T_{0}(1)^{-}$ to be the inverse image of $T_{0}(1)\times\{0\}$, $\{0\}\times T_{0}(1)$ by the map $\mathrm{Mess}$ respectively. Then we have
@@
The anti-holomorphic energy $E(F_{\pm})$ of the induced harmonic maps $F_{\pm}$ defines a finite valued functional over $T^{*}T_{0}(1)$. Moreover, we have
@@
Here $i_{\pm}:T_{0}(1)^{\pm}\to T^{*}T_{0}(1)$ denotes the inclusion map of $T_{0}(1)^{\pm}$ respectively.
Combining (1.1), (1.2), Theorem 1.1, and Theorem 1.2, we have
@@
The following equality holds over $T_{0}(1)^{\pm}$,
@@
Corollary 1.3 states that the anti-holomorphic energy functional plays the same role as the Liouville action $S$ or equivalently the Loewner energy $\pi I_{L}$ up to the symplectic diffeomorphism $\mathrm{Mess}$.
Here is an explanation of structure of this paper. In Section 2, we introduce some basic materials on the universal Teichmüller space $T(1)$ and the Weil-Petersson universal Teichmüller space $T_{0}(1)$. In Section 3, we introduce the definition of the maximal disc of Weil-Petersson class and study some properties of it. In Section 4 and 5, we prove main theorems using the variational methods which have been used in [16], [17], [14], [15], [12], and [11]. In the appendix, we present some basics about the 3-dimensional Anti de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$.
@@ 2. Universal Teichmüller Space
This section is a brief introduction of the universal Teichmüller space and the Weil-Petersson universal Teichmüller space. We refer to the chapter 1 of [9] and the chapter 16 of [6] for the universal Teichmüller space and the chapter 1 of [15] for the Weil-Petersson universal Teichmüller space.
@@ 2.1. Universal Teichmüller space
Let $\mathrm{QS}(\mathbb{S}^{1})$ denote the group of the quasisymmetric homeomorphism of the circle $\mathbb{S}^{1}$. Then the universal Teichmüller space is defined by
@@
where $\mathrm{Mob}(\mathbb{S}^{1})\cong\mathrm{PSL}(2,\mathbb{R})$ denotes the subgroup preserving $\mathbb{S}^{1}$ of the Möbius transformation group $\mathrm{PSL}(2,\mathbb{C})$ acting on ${}^\widehat{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$.
Let $\mathbb{D}=\{\,z\in\mathbb{C}\,|\,|z|<1\,\}$ be the open unit disk and let $\mathbb{D}^{*}=\{\,z\in\mathbb{C}\,|\,|z|>1\,\}$ be its exterior in $\widehat{\mathbb{C}}$. Denote by $L^{\infty}(\mathbb{D})$ and $L^{\infty}(\mathbb{D}^{*})$ the complex Banach spaces of bounded Beltrami differentials on $\mathbb{D}$ and $\mathbb{D}^{*}$ respectively, and let $L^{\infty}(\mathbb{D})_{1}$ be the unit ball in $L^{\infty}(\mathbb{D})$. Given a Beltrami differential $\mu\in L^{\infty}(\mathbb{D})_{1}$, extend it to $\mathbb{D}^{*}$ by the reflection
@@
with fixed points $1,-1,-\sqrt{-1}$. Then $w_{\mu}$ preserves $\mathbb{S}^{1}$ and $w_{\mu}|_{\mathbb{S}^{1}}\in\mathrm{QS}(\mathbb{S}^{1})$. Conversely, by extension theorem of Beurling-Ahlfors, any quasisymmetric homeomorphism in $\mathrm{QS}(\mathbb{S}^{1})$ can be extended to a quasiconformal homeomorphism $w_{\mu}$ of $\mathbb{D}$ for a $\mu\in L^{\infty}(\mathbb{D})_{1}$, hence we have the following description of $T(1)$,
@@
where $\mu\sim\nu$ if and only if $w_{\mu}|_{\mathbb{S}^{1}}=w_{\nu}|_{\mathbb{S}^{1}}$. We denote the equivalence class of $\mu$ by $[\mu]\in T(1)$. The space $T(1)$ has a unique structure of a complex Banach manifold such that the projection map
@@
is a complex linear surjection of holomorphic tangent spaces. The kernel of $D_{0}\mathrm{P}$ is the subspace $\mathcal{N}(\mathbb{D})$ of infinitesimally trivial Beltrami differentials.
@@
with the normalization $w^{\mu}(0)=0$, $(w^{\mu})_{z}(0)=1$ and $(w^{\mu})_{zz}(0)=0$. Then $w^{\mu}$ is conformal on $\mathbb{D}^{*}$. Then, we have
@@
For every $\mu\in L^{\infty}(\mathbb{D})_{1}$, the holomorphic function $\mathcal{S}(w^{\mu}|_{\mathbb{D}^{*}})$ lies in $A_{\infty}(\mathbb{D}^{*})$ by Kraus-Nehari inequality. Let us define the Bers embedding by
$\beta([\mu])=\mathcal{S}(w^{\mu}|_{\mathbb{D}^{*}})\in A_{\infty}(\mathbb{D}^{*}).$
The Bers embedding is a holomorphic map of complex Banach manifolds.
@@
where $e^{\psi}$ denotes the hyperbolic density function on $\mathbb{D}$ and $A_{\infty}(\mathbb{D})$ is defined as $A_{\infty}(\mathbb{D}^{*})$. The decomposition
@@
identifies the holomorphic tangent space $T_{[0]}T(1)=L^{\infty}(\mathbb{D})/\mathcal{N}(\mathbb{D})$ at the origin in $T(1)$ as follows
@@
The complex linear mapping $D_{0}\beta$ induces the isomorphism $\Omega^{-1,1}(\mathbb{D})\cong A_{\infty}(\mathbb{D}^{*})$ of the holomorphic tangent spaces to $T(1)$ and $A_{\infty}(\mathbb{D}^{*})$ at the origin.
The unit ball $L^{\infty}(\mathbb{D})_{1}$ carries a group structure induced by the composition of quasiconformal maps. The group law $\lambda=\nu\star\mu^{-1}$ is defined through $w_{\lambda}=w_{\nu}\circ w_{\mu}^{-1}$. The explicit formula of the group law is given by
@@
For $\mu\in L^{\infty}(\mathbb{D})_{1}$, using the group structure of $L^{\infty}(\mathbb{D})_{1}$, we define the right translation $R_{\mu}$ on $L^{\infty}(\mathbb{D})_{1}$. The induced right translations on $T(1)$
@@
are biholomorphic automorphisms of $T(1)$. Therefore, $D_{0}R_{[\mu]}:T_{[0]}T(1)\to T_{[\mu]}T(1)$ is a complex linear isomorphism between holomorphic tangent spaces, and we obtain the identification of $T_{[\mu]}T(1)\cong\Omega^{-1,1}(\mathbb{D})$.
@@
are biholomorphic as functions on the Banach space $A_{\infty}(\mathbb{D}^{*})$. The structure of $T(1)$ as a complex Banach manifold modeled on the Banach space $A_{\infty}(\mathbb{D}^{*})$ is explicitly described by the complex-analytic atlas given by the open covering
@@
with coordinate maps $h_{\mu}$ and the transition maps $h_{\mu\nu}$. Complex coordinates on $T(1)$ defined by the coordinate charts $(U_{\mu},h_{\mu})$ are called Bers coordinates. For every $\nu\in\Omega^{-1,1}(\mathbb{D})$, set $\phi=D_{0}\beta(\nu)$ and define a holomorphic vector field $\frac{\partial}{\partial\epsilon_{\nu}}$ on $U_{0}$ by setting

### v2

@@ Abstract.
We introduce maximal discs of Weil-Petersson class in the 3-dimensional Anti-de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, whose parametrization space can be identified with the cotangent bundle $T^{*}T_{0}(1)$ of Weil-Petersson universal Teichmüller space $T_{0}(1)$. We prove that the Mess map defines a symplectic diffeomorphism from $T^{*}T_{0}(1)$ to $T_{0}(1)\times T_{0}(1)$, with respect to the canonical symplectic form on $T^{*}T_{0}(1)$ and the difference of pullbacks of the Weil-Petersson symplectic forms from each factor of $T_{0}(1)\times T_{0}(1)$. Furthermore, we show that the functional given by the anti-holomorphic energies of the induced Gauss maps associated with maximal discs of Weil-Petersson class serves as a Kähler potential for the restriction of the canonical symplectic form to certain submanifolds $T_{0}(1)^{\pm}\subset T^{*}T_{0}(1)$, which bijectively parametrize the space of maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$.
2020 Mathematics Subject Classification 30F60, 32G15, 53C50, 53A10
@@ 1. Introduction
In their seminal work [18], Takhtajan and Teo introduced a distinguished subspace of the universal Teichmüller space $T(1)$, endowed with a natural Hermitian structure. This subspace, denoted $T_{0}(1)$, is known as the Weil-Petersson universal Teichmüller space, as it carries a rich geometric structure induced by the Weil-Petersson inner product and its associated symplectic 2-form $\omega_{\mathrm{WP}}$. Moreover, in [18], they introduced the universal Liouville action $S$ for elements in $T_{0}(1)$, and proved that $S$ serves as a Kähler potential for $\omega_{\mathrm{WP}}$, satisfying the identity
(1.1) ¯$\partial\bar{\partial}S=-2\sqrt{-1}\,\omega_{\mathrm{WP}}\qquad\text{over}\quad T_{0}(1),$
where $\partial$ and $\bar{\partial}$ denote the holomorphic and anti-holomorphic derivatives on $T_{0}(1)$. The identity (1.1) can be regarded as a universal analogue of the results previously established for classical Liouville actions on Teichmüller spaces of Riemann surfaces, as presented in [19], [20], [17], [14], and [15].
An element of $T_{0}(1)$ can be represented by a quasi-circle in the complex plane satisfying a specific condition, and such curves are referred to as Weil-Petersson curves. Since their introduction in [18], these curves have attracted significant attention and have become a central object of study in various areas of mathematics. In particular, Wang, in [26], introduced the Loewner energy $I_{L}$ for Weil-Petersson curves, defined via the energy of the driving function naturally associated with the Shramm Loewner evolution, and proved the identity
@@
This identity is particularly striking, as it equates two quantities defined through entirely different frameworks - one arising from Teichmüller theory and the other from Schramm-Loewner theory.
In a recent work [3], Bishop undertook a detailed study of Weil-Petersson curves from both analytical and geometric perspectives. In particular, he introduced an invariant associated with a minimal surface in the hyperbolic 3-space $\mathbb{H}^{3}$ that bounds a given Weil-Petersson curve on the boundary ¯$\partial\overline{\mathbb{H}}^{3}$. However, a limitation of this approach is the non-uniqueness of such bounding minimal surfaces in $\mathbb{H}^{3}$, which complicates efforts to establish a direct connection between the invariant of the bounding minimal surface in $\mathbb{H}^{3}$ and either the universal Liouville action or, equivalently, the Loewner energy of the Weil-Petersson curve.
Motivated by the previously described developments in [18], [26], and [3], we consider maximal surfaces in the Anti de Sitter 3-space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, which can be viewed as Lorentzian analogues of minimal surfaces in $\mathbb{H}^{3}$. More precisely, we consider a maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, whose boundary at infinity is the graph of a quasisymmetric homeomorphism of $\mathbb{S}^{1}$ representing an element in $T_{0}(1)$. Such a bounding maximal disc exists uniquely for each element in $T_{0}(1)$ by the result of [5]. Given a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, we consider a conformal embedding
$\sigma:\mathbb{D}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$
such that $\sigma(\mathbb{D})=\Sigma$ where $\mathbb{D}$ is the unit disc in the complex plane. Then, associated with the Gauss map of this maximal conformal embedding into $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, there exists a pair of harmonic maps
$F_{\pm}:\mathbb{D}\to\mathbb{D}.$
Here harmonicity encodes specific geometric structures on $\mathbb{D}$, a detailed discussion of which will be given later. In general, the integrals over $\mathbb{D}$ of the anti-holomorphic energy densities of the harmonic maps $F_{\pm}$ diverge. However, in Proposition 3.8, we show that the anti-holomorphic energies of the harmonic maps $F_{\pm}$ are finite, provided that the maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ satisfies the Weil-Petersson condition. As we will state more precisely, the anti-holomorphic energy of $F_{\pm}$ plays the role of the Liouville action $S$, or equivalently the Loewner energy $I_{L}$.
To formulate this result, we require a geometric framework based on the symplectic geometry of the holomorphic cotangent bundle $T^{*}T(1)$ and so called $\mathrm{Mess}$ map. We observe that $T^{*}T(1)$ parametrizes all conformal embeddings into $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, as we will elaborate on in Section 3. In this context, we introduce the Mess map, following the original construction by Mess in [12]:
@@
In this paper, we prove that the restriction
$\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$
is a symplectic diffeomorphism with respect to the canonical symplectic form $\omega_{\mathrm{C}}$ on $T^{*}T_{0}(1)$ and the difference of pullbacks of Weil-Petersson symplectic forms from each factor of $T_{0}(1)\times T_{0}(1)$. Here, the canonical symplectic form $\omega_{C}$ is defined as the imaginary part of the complex canonical symplectic form $\omega_{\mathbb{C}}$. See (4.23) for its precise definition. With this background in place, our main result concerning the $\mathrm{Mess}$ map is stated as follows:
@@
The map
$\mathrm{Mess}:T^{*}T_{0}(1)\to T_{0}(1)\times T_{0}(1)$
is a symplectic diffeomorphism. That is,
$\omega_{\mathrm{C}}=-\mathrm{Mess}^{*}_{+}(\omega_{\mathrm{WP}})+\mathrm{Mess}^{*}_{-}(\omega_{\mathrm{WP}}),$
where $\mathrm{Mess}_{\pm}:=\pi_{\pm}\circ\mathrm{Mess}$, and $\pi_{\pm}$ denote the projection maps from $T_{0}(1)\times T_{0}(1)$ onto the first or second factors, respectively.
A related result to Theorem 1.1 was presented in [9], where a similar claim was made for the $\mathrm{Mess}$ map at the origins of $T^{*}T(1)$ and $T(1)\times T(1)$. While the statement in [9] was formulated in terms of the Weil-Petersson symplectic form on $T(1)$, the precise nature of such a structure on $T(1)$ remains to be clarified, especially given that the construction of the Weil-Petersson symplectic form $\omega_{\mathrm{WP}}$ on $T_{0}(1)$ relies essentially on the underlying Hilbert manifold structure of $T_{0}(1)$ in [18]. This same Hilbert manifold structure on $T_{0}(1)$ also plays a crucial role in the proof of Theorem 1.1. See also Remark 4.12 for a related implication of the Hilbert structure on a fiber of the tangent bundle $TT_{0}(1)$.
Let us denote by $T_{0}(1)^{+}$, $T_{0}(1)^{-}$ the submanifolds of the cotangent bundle $T^{*}T_{0}(1)$, defined as the inverse images of $T_{0}(1)\times\{0\}$, $\{0\}\times T_{0}(1)$, respectively, under the $\mathrm{Mess}$ map. These submanifolds are real symplectic submanifold of the complex manifold $T^{*}T_{0}(1)$, equipped with the restriction of $\omega_{\mathrm{C}}$. Each of these submanifolds can also be interpreted as the image of a (non-holomorphic) section of $T^{*}T_{0}(1)$. Consequently, we can endow $T_{0}(1)^{\pm}$ with a holomorphic structure by identifying them with $T_{0}(1)$ via these sections. This phenomenon appears similarly in the derivation of the equality (1.1) in [18], where the corresponding submanifold of $T^{*}T_{0}(1)$ arises from the (non-holomorphic) section defined by the Schwarzians of conformal welding factors of given quasisymmetric homeomorphisms in $T_{0}(1)$. Further discussions on these are provided in Remark 5.6. As mentioned in Remark 3.13, the submanifolds $T_{0}(1)^{\pm}$ also admit a geometric interpretation: they provide a bijective parametrization of the space of maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. We then obtain the following result:
@@
The anti-holomorphic energy $E(F_{\pm})$ of the induced harmonic maps $F_{\pm}$ defines a finite-valued functional on $T^{*}T_{0}(1)$. Moreover, the following identity holds:
@@
Here $i_{\pm}:T_{0}(1)^{\pm}\to T^{*}T_{0}(1)$ denote the inclusion maps.
Combining equations (1.1), (1.2) with Theorems 1.1 and 1.2, we obtain the following:
@@
The following identity holds over $T_{0}(1)^{\pm}$:
@@
Remark 1.4.
Theorem 1.2 states that $T_{0}(1)^{\pm}$ admits a Kähler structure, where the complex structure is inherited via its identification with $T_{0}(1)$, and the Kähler form is given by $i^{*}\omega_{\mathrm{C}}$, up to a constant. Corollary 1.3 further shows that this Kähler structure on $T_{0}(1)^{\pm}$ is symplectically equivalent to the Weil-Petersson structure on $T_{0}(1)$. The discrepancy between these Kähler structures stems from the intrinsic differences of the anti-holomorphic energy functional $E$ and the Liouville action $S$, or equivalently the Loewner energy $\pi I_{L}$. For a more detailed discussion on the relation between $E$ and $S=\pi I_{L}$, we refer the reader to Remark 5.10.
Here is an explanation of structure of this paper. In Section 2, we review foundational material on the universal Teichmüller space $T(1)$ and the Weil-Petersson universal Teichmüller space $T_{0}(1)$. Section 3 introduces the notion of the maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ and investigates their fundamental properties. Section 4 and 5 are devoted to the proofs of the main theorems, employing variational techniques which played crucial roles in [19], [20], [17], [18], [14], [15], and [13]. Finally, the appendix provides a brief overview of the 3-dimensional Anti de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$.
@@ 2. Universal Teichmüller Space
This section provides a brief introduction to the universal Teichmüller space and the Weil-Petersson universal Teichmüller space. For further details, we refer to Chapter 1 of [11] and Chapter 16 of [8] for the universal Teichmüller space, and Chapter 1 of [18] for the Weil-Petersson universal Teichmüller space.
@@ 2.1. Universal Teichmüller space
Let $\mathrm{QS}(\mathbb{S}^{1})$ denote the group of the quasisymmetric homeomorphism of the circle $\mathbb{S}^{1}$. The universal Teichmüller space is then defined by
@@
where $\mathrm{Mob}(\mathbb{S}^{1})\cong\mathrm{PSL}(2,\mathbb{R})$ is the subgroup of the Möbius transformation group $\mathrm{PSL}(2,\mathbb{C})$ that preserves $\mathbb{S}^{1}$ and acts on ${}^\widehat{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$.
Let $\mathbb{D}=\{\,z\in\mathbb{C}\,|\,|z|<1\,\}$ denote the open unit disk and let $\mathbb{D}^{*}=\{\,z\in\mathbb{C}\,|\,|z|>1\,\}$ be its exterior in $\widehat{\mathbb{C}}$. Denote by $L^{\infty}(\mathbb{D})$ and $L^{\infty}(\mathbb{D}^{*})$ the complex Banach spaces of bounded Beltrami differentials on $\mathbb{D}$ and $\mathbb{D}^{*}$ respectively. Let $L^{\infty}(\mathbb{D})_{1}$ denote the unit ball in $L^{\infty}(\mathbb{D})$. For a given Beltrami differential $\mu\in L^{\infty}(\mathbb{D})_{1}$, we extend it to $\mathbb{D}^{*}$ by the reflection
@@
with fixed points $1,-1,-\sqrt{-1}$. Then $w_{\mu}$ preserves $\mathbb{S}^{1}$ and satisfies $w_{\mu}|_{\mathbb{S}^{1}}\in\mathrm{QS}(\mathbb{S}^{1})$. Conversely, by extension theorem of Beurling-Ahlfors, any quasisymmetric homeomorphism in $\mathrm{QS}(\mathbb{S}^{1})$ can be extended to a quasiconformal homeomorphism $w_{\mu}$ of $\mathbb{D}$ for some $\mu\in L^{\infty}(\mathbb{D})_{1}$. This leads to the following description of the universal Teichmüller space:,
@@
where $\mu\sim\nu$ if and only if $w_{\mu}|_{\mathbb{S}^{1}}=w_{\nu}|_{\mathbb{S}^{1}}$. We denote the equivalence class of $\mu$ by $[\mu]\in T(1)$. The space $T(1)$ admits a unique structure of a complex Banach manifold such that the projection map
@@
is a complex linear surjection onto the holomorphic tangent space of $T(1)$. The kernel of $D_{0}\mathrm{P}$ is the subspace $\mathcal{N}(\mathbb{D})$ of infinitesimally trivial Beltrami differentials.
@@
with the normalization $w^{\mu}(0)=0$, $(w^{\mu})_{z}(0)=1$ and $(w^{\mu})_{zz}(0)=0$. Then $w^{\mu}$ is conformal on $\mathbb{D}^{*}$. This leads to the following characterization of the universal Teichmüller space:
@@
For every $\mu\in L^{\infty}(\mathbb{D})_{1}$, the holomorphic function $\mathcal{S}(w^{\mu}|_{\mathbb{D}^{*}})$ belongs to $A_{\infty}(\mathbb{D}^{*})$ by Kraus-Nehari inequality. The Bers embedding is then defined by
(2.7) $\beta([\mu])=\mathcal{S}(w^{\mu}|_{\mathbb{D}^{*}})\in A_{\infty}(\mathbb{D}^{*}).$
This embedding is a holomorphic map between complex Banach manifolds.
@@
where $e^{\psi}$ denotes the hyperbolic density function on $\mathbb{D}$ and $A_{\infty}(\mathbb{D})$ is defined analogously to $A_{\infty}(\mathbb{D}^{*})$. The decomposition
@@
identifies the holomorphic tangent space $T_{[0]}T(1)\cong L^{\infty}(\mathbb{D})/\mathcal{N}(\mathbb{D})$ at the origin in $T(1)$ as
@@
The complex linear mapping $D_{0}\beta$ induces an isomorphism $\Omega^{-1,1}(\mathbb{D})\cong A_{\infty}(\mathbb{D}^{*})$ between the holomorphic tangent spaces to $T(1)$ and $A_{\infty}(\mathbb{D}^{*})$ at the origin.
The unit ball $L^{\infty}(\mathbb{D})_{1}$ carries a group structure induced by the composition of quasiconformal maps. The group law $\lambda=\nu\star\mu^{-1}$ is defined via
$w_{\lambda}=w_{\nu}\circ w_{\mu}^{-1}.$
The explicit formula of the group law is given by
@@
For $\mu\in L^{\infty}(\mathbb{D})_{1}$, using this group structure, we define the right translation $R_{\mu}$ on $L^{\infty}(\mathbb{D})_{1}$. The induced right translations on $T(1)$
@@
are biholomorphic automorphisms of $T(1)$. Consequently, the differential
$D_{0}R_{[\mu]}:T_{[0]}T(1)\to T_{[\mu]}T(1)$
is a complex linear isomorphism between the holomorphic tangent spaces, leading to the identification $T_{[\mu]}T(1)\cong\Omega^{-1,1}(\mathbb{D})$.
@@
are biholomorphic as maps on the Banach space $A_{\infty}(\mathbb{D}^{*})$. The structure of $T(1)$ as a complex Banach manifold, modeled on the Banach space $A_{\infty}(\mathbb{D}^{*})$, is explicitly described by the complex-analytic atlas given by the open covering
@@
with coordinate maps $h_{\mu}$ and the transition maps $h_{\mu\nu}$. Complex coordinates on $T(1)$, defined by the coordinate charts $(U_{\mu},h_{\mu})$, are referred to as Bers coordinates. For every $\nu\in\Omega^{-1,1}(\mathbb{D})$, let $\phi=D_{0}\beta(\nu)$ and define a holomorphic vector field $\frac{\partial}{\partial\epsilon_{\nu}}$ on $U_{0}$ by setting